A man is three times as old as his son. Four years ago he was four times as old as his son was at that time. How old is the son
Answer by ankor@dixie-net.com(22740)


m = man's age
s = son's age
:
A man is three times as old as his son.
m = 3s
Four years ago he was four times as old as his son was at that time.
m - 4 = 4(s - 4)
m - 4 = 4s - 16
m = 4s - 16 + 4
m = 4s - 12
:
How old is the son
replace m with 3s
3s = 4s - 12
12 = 4s - 3s
s = 12 yrs old is the son
:
:
Check solution in the statement:
m = 3(12) = 36
"Four years ago he was four times as old as his son was at that time. "
36 - 4 = 4(12 - 4)
32 = 4(8)